### v1

@@ 2.1 Lemmas for $n\leq 5$ and $n=6$
The two special cases $n\leq 5$ and $n=6$ follow from Lemma 1 and 3. ∎
@@ 3 The Second-Order Cone (SOC)
In this section, we will let $T_{n}$ be the conical semigroup $\text{SOC}(n)\cap\mathbb{Z}^{n}$ where
${\small\text{SOC}(n)\mathrel{\mathop{\ordinarycolon}}=\left\{\mathbf{x}\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}0\leq\sqrt{x_{1}^{2}+\dots+x_{n-1}^{2}}\leq x_{n}\right\}.}$
Additionally, for $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$, consider the bilinear form
$\langle\mathbf{a},\mathbf{b}\rangle\mathrel{\mathop{\ordinarycolon}}=a_{1}b_{1}+a_{2}b_{2}+\dots+a_{n-1}b_{n-1}-a_{n}b_{n}.$
@@
Note that $G$ maps $\partial T_{n}$ to $\partial T_{n}$.
We will begin this section by discussing the structural properties of the sporadic points of $T_{n}$. Then we use the structural properties of Pythagorean tuples and sporadic points to prove Theorem 1.2.
@@ 3.1 Sporadic Points of $\text{SOC}(n)\cap\mathbb{Z}^{n}$
Next, we will provide some Lemmas about the properties of sporadic points necessary to prove Theorem 1.2. For a detailed exposition of the technical proofs, please refer to the extended version. Lemmas 6 and 7 show that sporadic points are close to the boundary, $\partial T_{n}$.
Lemma 6
Suppose $\mathbf{s}\in T_{n}$ is a primitive sporadic with nonnegative entries such that $s_{n}>1$ and $s_{i}\neq 0$ for some $i\in[n-1]$. Then,
@@
Lemma 7
Let $\mathbf{s}\in T_{n}$. If $\langle\mathbf{s},\mathbf{s}\rangle=-1$, then $\mathbf{s}$ is sporadic. When $73\leq n<$ 7, we also have the converse: if $\mathbf{s}$ is sporadic, then $\langle\mathbf{s},\mathbf{s}\rangle=-1$.
@@
3.2 Proof of Theorem 1.2
We now present a complete formulation of Theorem 1.2 followed by its proof.
Theorem 3.1

### v2

@@ 2.2 Proof of Theorem 1
The two special cases $n\leq 5$ and $n=6$ follow from Lemma 5 and 13. ∎
@@ 3 The Second-Order Cone (SOC)
In this section, we use $T_{n}$ to denote the conical semigroup $\text{SOC}(n)\cap\mathbb{Z}^{n}$ where
$\text{SOC}(n):=\left\{\mathbf{x}\in\mathbb{R}^{n}:0\leq\sqrt{x_{1}^{2}+\dots+x_{n-1}^{2}}\leq x_{n}\right\}.$
The last component $x_{n}$ of $\mathbf{x}\in T_{n}$ is referred to as the height of $\mathbf{x}$. Additionally, for $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$, consider the bilinear form
$\langle\mathbf{a},\mathbf{b}\rangle:=a_{1}b_{1}+a_{2}b_{2}+\dots+a_{n-1}b_{n-1}-a_{n}b_{n}.$
@@
We remark that any action in $G$ in the above lemma maps $\partial T_{n}$ to $\partial T_{n}$.
We will begin this section by discussing some structures of the sporadic points of $T_{n}$. Then we use these structures of Pythagorean tuples and sporadic points to prove Theorem 2.
@@ 3.1 Sporadic Points of $\text{SOC}(n)\cap\mathbb{Z}^{n}$
Proof.
The first claim follows by simply checking the required inequalities directly. For the second claim, we proceed by contradiction. Suppose $\mathbf{s}$ is sporadic but $(A_{n}^{+})\mathbf{s}$ is not. Then, there is some point $\mathbf{p}\in T_{n}$ such that $\langle\mathbf{p},\mathbf{p}\rangle=0$ and $(A_{n}^{+})\mathbf{s}-\mathbf{p}\in T_{n}$. However, we would then have that
$(A_{n}^{+})^{-1}(A_{n}^{+}\mathbf{s}-\mathbf{p})=\mathbf{s}-(A_{n}^{+})^{-1}\mathbf{p}\in T_{n}.$
As $\mathbf{p}\in\partial T_{n}$, $(A_{n}^{+})^{-1}\mathbf{p}\in\partial T_{n}$, which is a contradiction with $\mathbf{s}$ being sporadic. The case of $(A_{n}^{+})^{-1}\mathbf{s}$ being sporadic follows the same argument. ∎
Next, we will provide some lemmas about sporadic points that are necessary to prove Theorem 2. Lemmas 16 and 17 show that sporadic points are close to the boundary $\partial T_{n}$.
Lemma 16.
Suppose $\mathbf{s}\in T_{n}$ is a primitive sporadic with nonnegative entries such that $s_{n}>1$ and $s_{i}\neq 0$ for some $i\in\{1,\dots,n-1\}$. Then,
@@
Proof.
We will show this by proving that
$\sqrt{s_{1}^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}<s_{n}<\sqrt{s_{1}^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}+1$
where the first inequality is given by membership in $T_{n}$. Without loss of generality, we can assume that $s_{1}\neq 0$. By way of contradiction, suppose that $\mathbf{s}$ is a primitive sporadic such that $s_{n}>1$ and $s_{1}>0$, and that $s_{n}\geq\sqrt{s_{1}^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}+1$. Then, we would have that
$s_{n}-1\geq\sqrt{s_{1}^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}>\sqrt{(s_{1}-1)^{2}+s_{2}^{2}+\dots+s_{n-1}^{2}}$
which is equivalent to $\mathbf{s}-(1,0,\dots,0,1)\in T_{n}$. This contradicts the assumption that $s$ is sporadic. Thus, we have have the desired equality. ∎
Lemma 17.
Let $\mathbf{s}\in T_{n}$. If $\langle\mathbf{s},\mathbf{s}\rangle=-1$, then $\mathbf{s}$ is sporadic.
Proof.
By way of contradiction, suppose $\langle\mathbf{s},\mathbf{s}\rangle=-1$ and that $\mathbf{s}$ is not sporadic. Then, there exists some $\textbf{p}\in T_{n}$ such that $\langle\mathbf{p},\mathbf{p}\rangle=0$ and $\textbf{s}-\textbf{p}\in T_{n}$. This is equivalent to saying that
$\langle\mathbf{s}-\mathbf{p},\mathbf{s}-\mathbf{p}\rangle\leq 0.$
This gives us that
$\langle\mathbf{s},\mathbf{s}\rangle-2\langle\mathbf{s},\mathbf{p}\rangle+\langle\mathbf{p},\mathbf{p}\rangle=-1-2\langle\mathbf{s},\mathbf{p}\rangle$ $\leq 0.$
Thus, $\langle\mathbf{s},\mathbf{p}\rangle\geq-\frac{1}{2}$ so $\langle\mathbf{s},\mathbf{p}\rangle\geq 0$ by its integrality. However, as $\langle\mathbf{s},\mathbf{s}\rangle=-1$ implies that $\sqrt{s_{1}^{2}+\dots+s_{n-1}^{2}}<s_{n}$ and $\langle\mathbf{p},\mathbf{p}\rangle=0$ implies that $\sqrt{p_{1}^{2}+\dots+p_{n-1}^{2}}=p_{n}$, using the Cauchy-Schwarz inequality, we have that
$s_{1}p_{1}+\dots+s_{n-1}p_{n-1}\leq\sqrt{(s_{1}^{2}+\dots+s_{n-1}^{2})(p_{1}^{2}+\dots+p_{n-1}^{2})}<s_{n}p_{n}.$
Thus, $\langle\mathbf{s},\mathbf{p}\rangle<0$, reaching a contradiction. Therefore, $\langle\mathbf{s},\mathbf{s}\rangle=-1$ implies that $\mathbf{s}$ is sporadic. ∎
@@
Proof.
This is equivalent to showing that these are the only such sporadic points such that $s_{1}+s_{2}+s_{3}=s_{n}$. As $\mathbf{s}$ is sporadic, $\mathbf{s}-(1,0,\dots,0,1)\notin T_{n}$. This is equivalent to saying that $(s_{n}-1)^{2}<(s_{1}-1)^{2}+s_{s}^{2}+\dots+s_{n-1}^{2}$ or
$0\displaystyle 2s_{1}s_{2}+2s_{2}s_{3}+2s_{1}s_{3}-2s_{2}-2s_{3}-s_{4}^{2}-\dots-s_{n-1}^{2}<$ 0 . (3)
We begin by showing that the first six coordinates must be equal. We proceed by contradiction in each of the below arguments.
  • Suppose that $s_{1}\geq s_{2}+1$. Then, (3) implies
    00 0 $>2s_{2}(s_{2}+1)+2s_{2}s_{3}+2(s_{2}+1)-2s_{2}-2s_{3}-s_{4}^{2}-\dots-s_{n-1}^{2}$
    $=2s_{2}^{2}+4s_{2}s_{3}-s_{4}^{2}-\dots-s_{n-1}^{2}\geq 0.$
    As this is a contradiction, we must have that $s_{1}=s_{2}$.
  • Suppose that $s_{2}\geq s_{3}+1$. Then, (3) implies
    00 0 $>2s_{1}(s_{3}+1)+2s_{3}(s_{3}+1)+2s_{1}s_{3}-2(s_{3}+1)-2s_{3}-s_{4}^{2}-\dots-s_{n-1}^{2}$
    $=4s_{1}s_{3}+2s_{3}^{2}-s_{4}^{2}-\dots-s_{n-1}^{2}+2s_{1}-2s_{3}-2$
    $\geq 2(s_{2}-s_{3}-1)\geq 0.$
    As this is a contradiction, we must have that $s_{1}=s_{2}=s_{3}$.
  • Suppose that $s_{3}\geq s_{4}+1$. Then, (3) implies
    00 0 $>6(s_{4}+1)^{2}-4(s_{4}+1)-s_{4}^{2}-\dots-s_{n-1}^{2}$
    $=5s_{4}^{2}-s_{5}^{2}-\dots-s_{n-1}^{2}+2s_{4}+2\geq 0$
    As this is a contradiction, we must have that $s_{1}=s_{2}=s_{3}=s_{4}$.
  • Suppose that $s_{4}\geq s_{5}+1$. Then, (3) implies
    00 0 $>5(s_{5}+1)^{2}-4(s_{5}+1)-s_{5}^{2}-\dots-s_{n-1}^{2}$
    $=4s_{5}^{2}-s_{6}^{2}-\dots-s_{n-1}^{2}+6s_{5}+1$
    $\geq 6s_{5}+1\geq 0.$
    As this is a contradiction, we must have that $s_{1}=s_{2}=s_{3}=s_{4}=s_{5}$.
  • Suppose that $s_{5}\geq s_{6}+1$. Then, (3) implies
    00 0 $>4(s_{6}+1)^{2}-4(s_{6}+1)-s_{6}^{2}-\dots-s_{9}^{2}$
    $=3s_{6}^{2}-s_{7}^{2}-\dots-s_{n-1}^{2}+4s_{6}-4\geq 0.$
    As this is a contradiction, we must have that $s_{1}=s_{2}=s_{3}=s_{4}=s_{5}=s_{6}$.
This implies that we have no such sporadic points for $n\leq 6$. Suppose $n=7$. Then, any candidate tuple must be of one of the following form:
$(k,k,k,k,k,k,3k)$
where $k\in\mathbb{Z}_{>0}$. As $\mathbf{s}$ is assumed to be primitive, $k=1$ and the only possible tuple is $(1,1,1,1,1,1,3)$.
Suppose $n=8$. Then, any candidate tuple must be of one of the following forms:
$(k,k,k,k,k,k,s_{7},3k)$
$(k,k,k,k,k,k,k,3k)$
where $k\in\mathbb{Z}_{>0}$ and $s_{7}\leq k-1$. As $\mathbf{s}$ is assumed to be primitive, the second possible form only contributes the tuple $(1,1,1,1,1,1,1,3)$. Suppose $\mathbf{s}$ is of the first form listed. We claim that $k=1$. By way of contradiction, suppose that $k\geq 2$. Then, $\mathbf{s}-(1,0,\dots,0,1)\in T_{n}$ as
$(3k-1)^{2}-(k-1)^{2}-5k^{2}-(k+1)=3k^{2}-5k+1\geq 3>0$
Thus, the only tuple satisfying these restrictions is $(1,1,1,1,1,1,0,3)$.
Suppose $n=9$. Then, for $k\in\mathbb{Z}_{>0}$, any candidate tuple must be of one of the following forms:
$(k,k,k,k,k,k,s_{7},s_{8},3k)$ (4)
$(k,k,k,k,k,k,k,s_{8},3k)$ (5)
$(k,k,k,k,k,k,k,k,3k)$ (6)
where $s_{7},s_{8}\leq k-1$.
Suppose $\mathbf{s}$ is of the form (4). By way of contradiction, suppose that $k\geq 2$. We claim that $\mathbf{s}-(1,0,\dots,0,1)\in T_{n}$. This follows from that fact that
$(3k-1)^{2}-(k-1)^{2}-5k^{2}-s_{7}^{2}-s_{8}^{2}$ $\geq(3k-1)^{2}-(k-1)^{2}-5k^{2}-2(k-1)^{2}$
$=k^{2}-2$
$\geq 2>0$
Thus, $\mathbf{s}$ must not be sporadic and $k=1$. This gives us the tuple $(1,1,1,1,1,1,0,0,3)$.
Suppose $\mathbf{s}$ is of the form (5). By way of contradiction, suppose $k\geq 3$. Then, we claim that $\mathbf{s}-(1,0,\dots,0,1)\in T_{n}$. This follows from the fact that
$(3k-1)^{2}-(k-1)^{2}-6k^{2}-s_{8}^{2}$ $\geq(3k-1)^{2}-(k-1)^{2}-6k^{2}-(k-1)^{2}$
$=2k^{2}-2k-1$
$\geq 2>0.$
Thus, we only need to consider $k=1,2$. If $k=1$, this gives us the tuple $(1,1,1,1,1,1,1,0,3)$. Suppose $k=2$. This gives us the following possible tuples:
$(2,2,2,2,2,2,2,0,6),\hskip 14.22636pt(2,2,2,2,2,2,2,1,6).$
This first tuple listed is not primitive so this only give us the tuple $(2,2,2,2,2,2,2,1,6)$. Lastly, if $\mathbf{s}$ is of the form (6), then $\mathbf{s}$ is only primitive if $k=1$. This gives us our last tuple, $(1,1,1,1,1,1,1,1,3)$.
Suppose $n=10$. Then,for $k\in\mathbb{Z}_{>0}$, any candidate tuple must be of one of the following forms:
$(k,k,k,k,k,k,s_{7},s_{8},s_{9},3k)$ (7)
$(k,k,k,k,k,k,k,s_{8},s_{9},3k)$ (8)
$(k,k,k,k,k,k,k,s_{9},3k)$ (9)
$(k,k,k,k,k,k,k,k,3k)$ (10)
where $s_{7},s_{8},s_{9}\leq k-1$. Any tuple of form (10) is a Pythagorean tuple so we may exclude it. Suppose $\mathbf{s}$ is of the form (7). By way of contradiction, suppose $k\geq 2$. Then, we claim that $\mathbf{s}-(1,0,\dots,0,1)\in T_{n}$. This follows from the fact that
$(3k-1)^{2}-$ $(k-1)^{2}-5k^{2}-s_{7}^{2}-s_{8}^{2}-s_{9}^{2}$
$\geq(3k-1)^{2}-(k-1)^{2}-5k^{2}-3(k-1)^{3}$
$=2k-3\geq 1>0.$
Thus, $k=1$ and the only tuple we have of this form is $(1,1,1,1,1,1,0,0,0,3)$.
Suppose $\mathbf{s}$ is of the form (8). By way of contradiction, suppose $k\geq 2$. If $s_{9}\geq 1$, then $\mathbf{s}-(1,1,\dots,1,3)\in T_{n}$ as
$(3k-3)^{2}-7(k-1)^{2}-(s_{8}-1)^{2}-(s_{9}-1)^{2}$ $\geq(3k-3)^{2}-7(k-1)^{2}-2(k-2)^{2}$
$=4k-6\geq 2>0$
Thus $s_{9}=0$. Suppose $k\geq 3$. Then, $\mathbf{s}-(1,0,\dots,0,1)\in T_{n}$ as
$(3k-3)^{2}-(k-1)^{2}-6k^{2}-s_{8}^{2}$ $\geq(3k-3)^{2}-(k-1)^{2}-6k^{2}-(k-1)^{2}$
$=k^{2}-2k-1\geq 2>0$
Thus, our options are $k=1,2$. If $k=1$, this recovers the tuple $(1,1,1,1,1,1,1,0,0,3)$. If $k=2$, this gives us potential tuples $(2,2,2,2,2,2,2,0,0,6)$ and $(2,2,2,2,2,2,2,1,0,6)$. The first is not primitive so we exclude it.
Lastly, suppose $\mathbf{s}$ is of the form (9). If $s_{9}\neq 0$, the $\mathbf{s}-(1,1,\dots,1,3)\in T_{n}$. This follows from the fact that
$(3k-3)^{2}-8(k-1)^{2}-(s_{9}-1)^{2}$ $\geq(3k-3)^{2}-8(k-1)^{2}-(k-2)^{2}$
$=14k-11\geq 0$
Thus, $s_{9}=0$. The only primitive sporadic satisfying these constraints is $(1,1,1,1,1,1,1,1,0,3)$. This completes the proof. ∎
@@
3.2 Proof of Theorem 2
We now present a complete formulation of Theorem 2 followed by its proof.
Theorem 20.